The vector is given by R=3i+4j-5k find the magnitude of the xy component and the angle between R and xy axis

1
Expert's answer
2022-05-28T08:22:59-0400

The magnitude of the xy component is given by

"R_{xy}=\sqrt{3^2+4^2}=5"

The angle between R and xy axis

"\theta=\cos^{-1}\frac{R_{xy}}{R}"

"\theta=\cos^{-1}\frac{5}{5\sqrt{2}}=45^\circ"